OpenStudy (anonymous):

A 24-foot beam is to be cut in two. If one portion os half the length of the other portion, what are the two lengths of the portions?

OpenStudy (anonymous):

portion 1 = 1/2 * portion 2. Portion 1 + portion 2 = 24ft portion 1 = 1/2* (24 - portion 1) portion 1 = 12 * 2/3 = 8 portion 2 = 16

OpenStudy (anonymous):

awesome! thanks

OpenStudy (anonymous):

Cool.

OpenStudy (anonymous):

p + 2p = 24 ft -> p = 24/3 = 8 ft Thus 2p = 2 * 8 = 16 ft